OpenStudy (anonymous):

Find the slope-intercept (y = mx + b) form of the line satisfying the given conditions. Parallel to y = -2x, passing through (0,1)

OpenStudy (sburchette):

In order for the line to be parallel to y=-2x, it must have the same slope. The point (0,1) happens to be the y-intercept. So from this information, we know m and b. The equation of the line is\[y=-2x+1\]

OpenStudy (anonymous):

y-y1=m(x-x1) m=0, y1=1 and x1=0 y-1=-2(x-0) y-1=-2x Slope-intercept form is y=mx+c y=-2x+1

OpenStudy (anonymous):

ummmm i dont get it iam sorry

OpenStudy (anonymous):

Are you familiar with the equation \[m=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

And this is the same thing as \[y_2-y_1=m(x_2-x_1)\]

OpenStudy (anonymous):

And like @SBurchette was saying earlier parallel lines have the same slope. So the slope of a line parallel to y=-2x is -2 because m here is -2.

OpenStudy (anonymous):

i was working with that a little last week

OpenStudy (anonymous):

And it's okay if you don't understand but what exactly is it that you don't understand?

OpenStudy (anonymous):

Did you figure it out @compaqangel0205 if you want to start over we can do so in another question :)

OpenStudy (anonymous):

i was saying its all hard but it would not post ty